OpenStudy (anonymous):

Help Please! Subject: Hypotenuse Q. 2 of 5 @ali2x2 @ab854

OpenStudy (anonymous):

Ok, what's the question?

OpenStudy (anonymous):

The hypotenuse of a right triangle is 50 millimeters long. One leg of the right triangle is 30 millimeters long. What is the length of the other leg? 20 millimeters 30 millimeters 40 millimeters 60 millimeters

OpenStudy (anonymous):

Again we use pyragothean theorem (a^2 + b^2 = c^2), but this time we set it up as follows: 50^2 = 30^2 + b^2

OpenStudy (anonymous):

Then 50^2 - 30^2 = b^2 2500 - 900 = b^2 b^2 = 1600 b = \[\sqrt{1600}\] b = 40 centimeters